### v1

# NAMET: Robust Massive Model Editing via Noise-Aware Memory Optimization

Abstract
Model editing techniques are essential for efficiently updating knowledge in large language models (LLMs). However, the effectiveness of existing approaches degrades in massive editing scenarios, particularly when evaluated with practical metrics or in context-rich settings. We attribute these failures to embedding collisions among knowledge items, which undermine editing reliability at scale. To address this, we propose NAMET (Noise-aware Model Editing in Transformers), a simple yet effective method that introduces noise during memory extraction via a one-line modification to MEMIT. Extensive experiments across six LLMs and three datasets demonstrate that NAMET consistently outperforms existing methods when editing thousands of facts.
1 Introduction
Large language models (LLMs) are increasingly employed as search engines and chatbots, as they excel at retrieving knowledge to answer user queries [4, 31, 34, 3]. However, they are prone to spreading misinformation about frequently updated topics due to outdated training data [32, 11, 33]. To address this issue, retraining or fine-tuning models for partial knowledge updates is proposed [1, 28], albeit with prohibitively expensive overhead. In contrast, recent advances in model editing (ME) techniques [20, 6] enable massive editing of thousands of factual associations concurrently at minimal data and computational cost, thereby rendering real-time knowledge updates feasible.
Figure 1: (LEFT) Comparison of editing efficacy under both metrics on the CounterFact dataset, with and without prefixes. The editing performance is evaluated using MEMIT [20], PMET [15], and NAMET. The loose metric compares the next-token probability of the target and original objects, while the practical metric checks whether the target object is explicitly generated. (RIGHT) Distribution of edited knowledge representations extracted using MEMIT and NAMET.
Despite the success of existing massive ME techniques, we observe that their reported performance is often overestimated due to overly loose evaluation metrics. In particular, most prior works assess editing quality by checking whether the model is more likely to generate the following tokens as the target object than the original one, whereas neglecting to evaluate whether the model’s output is consistent with the target object [20, 6]. Therefore, we advocate a “practical metric”, which measures the proportion of cases in which the edited model retrieves the target object and explicitly generates related output. This metric provides a more accurate reflection of real-world usage, as will be shown in our evaluation setting (see Section 6). Considering the left of Figure 1, the loose metric does not require the edited model to explicitly generate the target object when editing "Jeep Commander is created by Honda." Under the loose metric, an edited model that outputs "Jeep Commander is created by the American company Jeep" may still be considered a successful edit, whereas the strict metric would classify it as a failure. Under such evaluation criteria, these methods fail to maintain performance in massive editing settings, as shown on the left of Figure 1. Moreover, existing approaches exhibit substantial degradation when edited knowledge is preceded by prefixes [17], which is common in realistic question-answering tasks [24, 26]. This lack of robustness to contextual variations undermines their practicality in real-world applications.
To achieve robust massive editing under practical settings, we propose NAMET (Noise-Aware Model Editing in Transformers), which outperforms existing approaches under stricter evaluation criteria and exhibits strong robustness to contextual variations. We first identify “embedding collisions”, which reflect the distinctiveness of hidden representations among edited facts, as a primary cause of performance degradation in both massive editing and context-rich scenarios. NAMET mitigates this issue by introducing noise during target memory optimization via a simple one-line modification to MEMIT. As shown on the right of Figure 1, the introduced noise reduces the occurrence of embedding collisions, thereby enabling a greater number of edits to be performed simultaneously.
In this paper, extensive experiments on six LLMs demonstrate that NAMET consistently outperforms existing methods under more rigorous scenarios across the CounterFact, ZsRE, and Wiki-recent datasets. Specifically, on the CounterFact dataset, NAMET achieves up to a 15% improvement in editing efficacy over existing methods, and maintains over 90% editing efficacy across all evaluated LLMs when editing 10,000 facts. Furthermore, NAMET consistently maintains high editing accuracy even when edited items are preceded by up to 200-token prefixes, highlighting its robustness in realistic, context-rich settings, including chatbots and long-context QA tasks.
2 Related Work
Model Editing. Existing ME techniques can be classified into auxiliary-based [10, 22, 38, 36, 21] and location-based methods [19, 20, 15]. Auxiliary-based ME techniques preserve the original parameters, and introduce additional information to edit knowledge. SERAC [22] requires extra memory to store new edits and learn to reason over them to manipulate the model’s output. Location-based methods directly modify model parameters to edit knowledge without requiring any additional information. These methods assume that factual associations are stored in the feed-forward networks (FFNs) of the LLMs [8, 7, 5]. Building on these, ROME [19] first gains insights on the specific location of the knowledge through causal analysis. It proceeds to directly modify critical MLP layers to update factual associations. MEMIT [20] builds upon ROME to enable massive editing of thousands of facts concurrently. AlphaEdit [6] focuses on sequential editing, aiming to preserve both previously edited knowledge and the general capabilities of the LLM during successive edits.
Massive Editing. In practical applications, ME techniques may aim to update a model with hundreds or even thousands of facts simultaneously in order to keep up with the constantly evolving knowledge [13, 9]. However, auxiliary-based methods are limited in scalability, typically supporting only a few edits at a time [22]. In contrast, location-based methods are more scalable for massive editing. MEMIT [20] scales to edit 10,000 facts concurrently, and PMET [15] further improves performance by incorporating attention layers when updating the parameters of the FFNs. Despite their effectiveness and scalability, these methods have been shown to be fragile when handling prefixes during evaluation, which is a common scenario in real-world applications [17, 35, 18]. Moreover, we observe that their performance in massive editing is overestimated due to the loose metric. In this work, we propose NAMET, which achieves superior performance in massive editing under practical evaluation metrics, while also exhibiting greater robustness against long prefixes.
3 Preliminary: Editing Memory in LLMs
Previous works have shown that a pre-trained LLM has memorized many factual associations [23, 12, 25, 27]. These stored facts could be edited by modifying the MLP layers within FFN modules, based on the assumption that knowledge is stored in them in the form of key-value pairs [8, 7].
Figure 2: Illustration of the model editing problem.
In Figure 2, the MLP layer $W_{out}^{l}$ within FFN associates keys $k_{t}^{l}(x)=\sigma(W^{l}_{in}\gamma(h^{l-1}_{t}(x)))$ with memories $m^{l}_{t}(x)$ for the fact $x$. Given the critical mediating role of MLP layers in storing facts, Meng et al. [19] shows that it is sufficient to update $W_{out}^{l}$ to edit stored facts. We then optimize $W_{out}^{l}$ (abbreviated as $W_{1}$) as follows:
$W_{1}\triangleq\underset{\hat{W}}{\arg\min}(\sum_{i=1}^{N_{t}}||\hat{W}k_{i}^{t}-m_{i}^{t}||^{2}+\sum_{j=1}^{N_{p}}||\hat{W}k_{j}^{p}-m_{j}^{p}||^{2})$ (1)
Here, $k_{i}^{t}$ and $k_{j}^{p}$ denote the encoded subject representations for individual target and preserved fact $i$ and $j$, respectively, while $m_{i}^{t}$ and $m_{j}^{p}$ represent their corresponding memory vectors. We stack the keys and memories of totally $N_{t}$ target knowledge into matrices as $K_{t}=[k_{1}^{t}\ |\ k_{2}^{t}\ |\ \dots\ |\ k_{N_{t}}^{t}]$ and $M_{t}=[m_{1}^{t}\ |\ m_{2}^{t}\ |\ \dots\ |\ m_{N_{p}}^{t}]$. Similarly, we construct $K_{p}$ and $M_{p}$ for $N_{p}$ preserved facts. The objective in Equation 1 can then be optimized by solving the normal equations [20]:
$(W_{0}+\Delta)\begin{bmatrix}K_{p}&K_{t}\end{bmatrix}$ $=\begin{bmatrix}M_{p}&M_{t}\end{bmatrix}$ (2)
$W_{0}K_{p}$ $=M_{p}$ (3)
where we expand $W_{1}$ into $W_{0}+\Delta$. $W_{0}$ denotes the original (unedited) parameters that associate preserved keys with their memory representations. The final update to $W_{out}^{l}$ can be computed by multiplying both sides of Equation 2 by $\begin{bmatrix}K_{p}&K_{t}\end{bmatrix}^{T}$, and subtracting Equation 3 from Equation 2[20]:
$\Delta(C_{p}+K_{t}K_{t}^{T})=RK_{t}^{T}$ (4)
where $R=M_{t}-W_{0}K_{t}$ denotes new relations’ residual with respect to the original weights, which can also be written as $[r_{1}^{t}\ |\ r_{2}^{t}\ |\ \ldots\ |\ r_{N_{t}}^{t}]$. Since the pretraining data of the original model is not accessible, we approximate $C_{p}$ using a set of randomly sampled inputs from public datasets:
$C_{p}=\lambda E_{k^{p}}[k_{i}^{p}(k_{i}^{p})^{T}]$ (5)
The scalar $\lambda$ balances the influence between newly edited facts and preserved knowledge.
Despite their effectiveness, existing methods struggle when concurrently editing a large number of facts or when the edited knowledge is preceded by prefixes. Here, we termed $r_{i}k_{i}^{T}$ the residual-key pair, and $k_{i}k_{i}^{T}$ the key-key pair for target fact $i$ [^1] We omit superscript $t$ for clarity. We then analyze these challenges through the lens of embedding collision of residual-key (abbreviated as r-k) and key-key (abbreviated as k-k) pairs.
4 Motivation
In this section, we investigate the underlying causes of the challenges associated with massive editing, as illustrated in Figure 1. Specifically, we examine the decline in editing performance as the number of edited facts increases, and the specific failure cases where edits are preceded by prefixes. Our theoretical and empirical analyses suggest these issues stem from embedding collisions, reflecting the insufficient distinctiveness between r-k and k-k pairs in the embedding space.
4.1 Theoretical Analysis of Embedding Collision
We observe that by expanding $K_{t}$ and $R$ in Equation 4, the update equation can be reformulated as:
$\Delta\left(C_{p}+\sum_{i=1}^{N_{t}}k_{i}k_{i}^{T}\right)=\sum_{i=1}^{N_{t}}r_{i}k_{i}^{T}$ (6)
where the update $\Delta$ is determined by the aggregated r-k and k-k embeddings across all edited facts. As the number of edits increases, solving Equation 6 is more likely to cause embedding collisions between the r-k and k-k embeddings associated with a specific fact with other facts. This eventually leads to information interference and degraded editing performance.
To formalize the concept of embedding collision, we define two key requirements for the desired update $\Delta$: (1) The update should preserve the existing knowledge, expressed as $\Delta C_{p}=0$. (2) The update should ensure lossless reconstruction for each individual fact, formulated as $\Delta k_{i}k_{i}^{T}=r_{i}k_{i}^{T}$, where $\Delta$ is computed while considering all target facts. For notational simplicity, we define $S_{i}=k_{i}k_{i}^{T}$ and $T_{i}=r_{i}k_{i}^{T}$. Incorporating (1), an ideal $\Delta$ that meets (2) implies:
$\Delta\sum_{i=1}^{N_{t}}S_{i}=\sum_{i=1}^{N_{t}}T_{i}\quad\rightarrow\quad\Delta S_{i}=T_{i}\quad\text{for }i=1,2,\ldots,N_{t}$ (7)
However, the validity of Equation 7 is intuitively affected by the degree of orthogonality among the embeddings of different facts. We then define this orthogonality as embedding collision:
Definition 1 (Embedding Collision). Given two knowledge items $i$ and $j$, and an embedding matrix $E\in\{S,T\}$, the embedding collision between $i$ and $j$ is the orthogonality between their respective embedding vectors $E_{i}$ and $E_{j}$, which could be measured by their cosine similarity $cos(E_{i},E_{j})$.
We now formalize the connection between embedding collision and the editing performance of a specific knowledge item $i$ under massive editing. Specifically, we quantify the degree to which Equation 7 is established by analyzing the reconstruction loss $RL_{i}=\Delta S_{i}-T_{i}$ for each knowledge item. This relationship is formalized in the following theorem:
Theorem 1. The reconstruction loss for the $i$-th knowledge item is given by:
$RL_{i}=(T_{i}S^{+}S_{i}-T_{i})+\sum_{j\neq i}T_{j}S^{+}S_{i}$ (8)
where the first term, $T_{i}S^{+}S_{i}-T_{i}$, is amplified by the cosine similarity (collision) between $S_{i}$ and other $S_{j}$ embeddings for $j\neq i$. The second term, $\sum_{j\neq i}T_{j}S^{+}S_{i}$, is jointly amplified by the collision between $T_{i}$ and $T_{j}$, as well as between $S_{i}$ and $S_{j}$ for $j\neq i$.
This result demonstrates how embedding collisions impact the editing performance of individual knowledge items under massive editing. Specifically, stronger collision among knowledge items leads to increased individual reconstruction loss, ultimately reducing the overall effectiveness of massive editing. The complete proof is provided in Appendix A.
4.2 Empirical Study of Embedding Collision
Motivated by the above analysis, we thus hypothesize that fostering more “isolated” r-k and k-k representations across knowledge items would enhance capacity for massive concurrent editing. Such isolation is also vital for robustness against long prefixes, where collisions typically exacerbate performance degradation. This intuition motivates two hypotheses below, which we propose and empirically verify, to explain failures in existing ME techniques, followed by our introduction of a simple yet effective method to address these challenges.
Hypothesis 1. As the number of edited knowledge items increases, the r-k and k-k pairs associated with a specific item are more likely to collide with those of other items in the embedding space, leading to degraded editing performance for that specific item.
Hypothesis 2. Long prefixes exacerbate embedding collision issues under massive editing, leading to degraded editing performance when edited facts are evaluated with descriptive prefixes.
We proceed to verify the above hypotheses through carefully designed experiments.
[…]
LLaMA2-7B 200 98.5 0.09 66.66% 0.19 66.66% 21.79 100.00% 100.00%
500 90.0 0.20 74.00% 0.41 78.00% 21.21 74.00% 92.00%
1000 86.8 0.40 69.70% 0.72 72.73% 20.16 81.82% 95.45%
Deepseek-7B 200 99.5 0.11 100.00% 0.21 100.00% 4.35 100.00% 100.00%
500 98.6 0.27 85.71% 0.60 100.00% 4.96 100.00% 100.00%
1000 97.8 0.52 81.81% 1.03 100.00% 7.03 86.36% 100.00%
Empirical Verification of Hypothesis 1. We separately modify 200, 500, and 1,000 knowledge items in LLaMA2-7B [30] and Deepseek-7B [3] using MEMIT [20]. These facts are sampled from the CounterFact dataset [19]. We evaluate the editing accuracy of these items where no prefix precedes the edited query. We measure the embedding collision using the collision score, which is defined as follows:
Definition 2 (Collision Score). Given a knowledge item $i$, and an embedding matrix $E\in\{S,T,\lambda\}$, its collision score $V^{E}_{i}$ is computed as the average cosine similarity between its embedding and those of all edited facts:
$V^{E}_{i}=\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\cos(E_{i},E_{j})$ (9)
For each knowledge item, we compute the collision score for their r-k ($T$) and k-k ($S$) embeddings. To incorporate the influence of $C_{p}$, we also calculate the collision score of the k-k-$\lambda$ pairs, defined as $\frac{1}{n}C_{p}+k_{i}k_{i}^{T}$ for each item. We aggregate the collision scores across all items to quantify the overall degree of collision. To further validate Hypothesis 1, we identify the items whose target objects fail to be retrieved during evaluation. We then examine whether their corresponding r-k, k-k, and k-k-$\lambda$ collision scores rank in the top 50% among all edited items.
As shown in Table 1, the decline in overall editing accuracy for both models is correlated with an increase in embedding collision when more facts are edited. Specifically, for LLaMA2-7B, editing accuracy decreases from 98.5% to 86.8% as the number of edited facts rises from 200 to 1,000. Correspondingly, the collision scores increase from 0.09 to 0.40 for r-k pairs and from 0.19 to 0.72 for k-k pairs. At the individual level, editing performance is similarly impacted by collision degree. Approximately 70% of r-k and k-k pairs associated with failed edits rank within the top 50% of collision scores. Incorporating $C_{p}$ further amplifies this effect, raising the proportion of high-collision items among failure cases to 100% for 200 edited items and 81.82% for 1,000 items. Overall, more than 90% of all failed edits exhibit at least one of the three collision scores (r-k, k-k, or k-k-$\lambda$) ranking within the top 50% of the collision distribution.
We also observe that, despite exhibiting higher collision scores for both r-k and k-k pairs compared to LLaMA2-7B, Deepseek-7B still achieves superior editing accuracy. This discrepancy does not contradict our hypothesis, as comparisons between embeddings across models may be influenced by variations in inherent model settings. For a more precise comparison, we extract the principal components of r-k and k-k pairs from the same 100 items in both LLaMA2-7B and Deepseek-7B, and visualize them in Figure 3. The r-k and k-k embeddings in Deepseek-7B are significantly more dispersed than those in LLaMA2-7B, suggesting that Deepseek-7B’s embeddings are less prone to collisions. This observation aligns with its superior editing performance, further supporting our hypothesis.
Figure 3: PCA visualization of r-k and k-k embeddings for 100 edited facts at layer 3 in LLaMA2-7B and Deepseek-7B, using MEMIT.
Table 2: Impact of varying prefix lengths on editing performance.
Model Prefix Lens Editing Acc. low r-k Acc. top r-k Acc. low k-k Acc. top k-k Acc. low k-k-$\lambda$ Acc. top k-k-$\lambda$ Acc.
LLaMA2-7B 0 98.50% - - - - - -
5 84.15% 83.00% 47.00% 73.00% 39.00% 93.00% 27.00%
10 80.35% 79.00% 44.00% 68.00% 34.00% 90.00% 24.00%
50 77.40% 70.00% 37.00% 54.00% 28.00% 90.00% 21.00%
Deepseek-7B 0 99.50% - - - - - -
5 79.35% 82.00% 36.00% 88.00% 31.00% 71.00% 58.00%
10 84.25% 92.00% 54.00% 97.00% 38.00% 85.00% 58.00%
50 85.10% 91.00% 47.00% 90.00% 37.00% 91.00% 62.00%
Empirical Verification of Hypothesis 2. We proceed by editing 200 facts from the CounterFact dataset in LLaMA2-7B and Deepseek-7B using MEMIT. To evaluate each edit, we measure the average editing accuracy when the original factual association is preceded by prefixes of varying lengths. Specifically, we vary the prefix length from 5 to 50 tokens and report the average editing accuracy across all edited items. Additionally, we compute the average editing accuracy of the 10 items with the highest collision scores for r-k, k-k, and k-k-$\lambda$ pairs separately. We also evaluate the 10 items with the lowest collision scores for the same types of pairs, in order to illustrate how embedding collisions influence the robustness of edited knowledge under long-prefix conditions.
Table 2 demonstrates the degradation of editing performance when evaluating edited facts with prefixes. For LLaMA2-7B, the accuracy drops from 98.5% to 84.15% when preceded by 5 tokens, and further to 77.40% for 50 tokens. The degradation phenomenon is also observed for Deepseek-7B, where the editing accuracy drops around 15% when preceded by 50-token prefixes. The robustness to prefix perturbations varies significantly between items that are prone to embedding collisions and those that are not. Corresponding to Hypothesis 2, items with lower collision degrees exhibit editing accuracy comparable to the overall average, whereas items with higher collision degrees suffer a substantial decline, with accuracy dropping to approximately 30%. This discrepancy is especially pronounced when collision is measured using r-k pairs. For both models, items with low collision degrees achieve editing accuracy around 80%, exceeding that of collision-prone items by over 40%.
These results highlight connections between embedding collisions, and the effectiveness and robustness of massive editing. Since key embeddings are fixed by the model and residuals are derived from target memory embeddings, advanced ME methods can only improve the isolatedness of memory embeddings across different knowledge items.
5 Adjustment to Existing Model Editing Method and Metric
Figure 4: (LEFT) Illustration of changing padding strategy. (RIGHT) Overview and information flow of NAMET.
Motivated by the above results, we propose NAMET, an effective one-line modification to MEMIT, which generates more isolated memory embeddings through introducing noise during optimization. This improves both the model’s capacity to edit multiple facts concurrently under practical metric, and the robustness when evaluating the edited items with prefixes. We then elaborate on the details.
Mitigating Embedding Collisions via Noise Injection. Padding tokens, which carry no semantic content, are ideal for injecting noise during optimizing target memory embeddings. Based on this insight, we propose to change the padding strategy from right-padding to left-padding, thereby prepending [unk] tokens before factual prompts during memory optimization. Formally, let $h_{i}^{L}$ denote the hidden state at layer $L$ corresponding to the $i$-th target factual association (subject $s_{i}$, relation $r_{i}$, object $o_{i}$). We optimize upon it to obtain target hidden state $z_{i}$ via:
$z_{i}=h_{i}^{L}+\underset{\delta_{i}}{\arg\min}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log\mathds{P}_{G(h_{i}^{L}+=\delta_{i})}[o_{i}\mid f_{j}\oplus tp(s_{i},r_{i})]$ (10)
Here, $\delta_{i}$ is optimized to maximize the model’s confidence in predicting the target object $o_{i}$. $f_{j}\oplus tp(s_{i},r_{i})$ concatenates random prefixes $f_{j}$ with templated prompt $tp(s_{i},r_{i})$ for each target edit $i$. $G_{(h_{i}^{L}+=\delta_{i})}$ denotes that we add change $\delta_{i}$ to the hidden state $h_{i}^{L}$ during the model execution.
As illustrated in Figure 4, each edit is augmented with $N_{FP}-1$ prefixes of fixed length, resulting in $N_{FP}$ factual prompts per edit. The first prompt contains no prefix and is thus shorter, causing the tokenizer to prepend [unk] tokens under left-padding. These padding tokens introduce noise in hidden states, promoting more isolated memory representations during optimization and mitigating embedding collisions across knowledge items. Notably, this method requires only a one-line modification to the original MEMIT implementation. The complete algorithm is provided in Appendix C.
Practical Evaluation Metric. We propose to evaluate the editing performance of ME techniques by requiring the edited models to strictly examine whether explicit target objects are retrieved, as demonstrated in Figure 1. The editing efficacy is then defined as:
$\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ (s_{i},r_{i}))]$ (11)
6 Experiments
In this section, we empirically focus on evaluating the following research questions (RQs), while additional experiments are provided in Appendix D. Codes are available at https://github.com/ybdai7/NAMET-massive-editing.
  • RQ1. Can NAMET generate more isolated embeddings for different knowledge items?
  • RQ2. How does NAMET perform on massive editing tasks compared with baselines for various LLMs? Can it excel under the practical metric?
  • RQ3. How does NAMET perform when evaluating edited facts with prefixes?
6.1 Experiments Setup
Models, datasets, and baselines. We conduct extensive experiments on various LLMs, including LLaMA2-7B [30], LLaMA2-13B [30], Falcon-7B [2], Qwen-2.5-7B [34], Deepseek-base-7B [3], and LLaMA3-8B [31]. We consider a range of ME techniques as baselines: FT [39], MEND [21], ROME [19], MEMIT [20], PMET [15], and ALPHAEDIT [6]. We demonstrate their performance on CounterFact [19], ZsRE [14], and Wiki-recent [37]. We provide a full description in Appendix B.1.
Evaluation metrics. Following previous work, we evaluate the performance of ME techniques in terms of efficacy, generalization, specificity, and fluency for CounterFact and ZsRE datasets. For Wiki-recent, we additionally evaluate the portability [37] of edited models, which represents the ability to address downstream tasks with edited knowledge. We evaluate the editing efficacy and generalization using the practical metric in Equation 11. When evaluating efficacy on the CounterFact and Wiki-recent datasets, and generalization on CounterFact, we prepend each prompt with 10 distinct 5-token prefixes. Full details of metrics are provided in Section B.2.
Figure 5: PCA visualization of r-k embeddings for 100 edited facts across different models.
6.2 Distinctiveness of Retrieved Embeddings (RQ1)
Finding 1. NAMET promotes more isolated r-k embeddings. We visualize the r-k pairs of 100 facts edited by MEMIT, PMET, and NAMET at layer 4 across models. As shown in Figure 5, the r-k pairs generated by NAMET are more dispersed in the embedding space, while those produced by MEMIT and PMET are more likely to form denser clusters. This observation supports our hypothesis that NAMET encourages more isolated and less entangled target memory embeddings.
6.3 Performance of Massive Editing (RQ2)
Table 3: Performance comparison of different editing methods on six LLMs over the Counterfact, Wiki-recent, and ZsRE benchmarks. We report the average value calculated over five evaluations.
Model Method Counterfact Wiki-recent ZsRE
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Flu.$\uparrow$ Eff.$\uparrow$ Por.$\uparrow$ Loc.$\uparrow$ Flu.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$
LLaMA2-7B FT 0.29 0.23 77.43 4.90 7.23 41.61 36.52 4.91 5.30 4.31 14.69
MEND 0.23 0.31 78.55 3.07 0.00 34.67 37.46 2.69 0.00 0.00 0.50
ROME 0.00 0.00 50.73 4.67 76.73 49.31 51.51 4.97 37.29 6.86 10.27
MEMIT 24.95 22.68 63.84 5.06 34.75 44.93 46.72 5.04 76.63 64.06 15.57
PMET 74.22 46.45 72.47 5.07 81.84 51.11 53.16 4.97 77.29 71.40 16.54
ALPHAEDIT 0.51 0.53 51.14 5.01 0.07 35.34 37.48 5.27 44.26 35.83 12.65
NAMET 89.09 61.21 69.19 5.15 82.51 53.13 54.61 5.08 77.88 73.42 16.65
Qwen2.5-7B FT 16.18 14.15 56.07 5.27 21.17 51.40 51.50 5.15 14.30 13.00 39.28
MEND 0.01 0.06 70.73 2.82 0.00 42.55 44.37 2.72 0.00 0.00 0.09
ROME 0.00 0.00 49.83 5.23 16.28 46.52 46.61 5.02 4.10 3.43 1.30
MEMIT 90.06 63.86 70.53 5.29 94.88 56.97 61.23 5.10 54.12 42.96 31.57
PMET 65.71 52.84 63.14 5.18 82.39 58.38 57.59 5.11 53.58 46.59 36.50
ALPHAEDIT 83.15 55.70 67.16 5.14 94.16 57.17 59.45 5.10 44.52 34.98 25.52
NAMET 90.49 64.37 72.18 5.36 95.61 57.46 60.28 5.09 66.29 60.22 41.80
LLaMA2-13B FT 1.23 0.07 68.57 4.84 13.90 36.89 40.09 4.97 5.95 5.10 16.16
ROME 4.05 1.52 50.44 5.25 11.06 38.14 39.09 4.47 5.52 5.06 2.25
MEMIT 47.98 34.75 71.61 5.17 94.76 51.38 50.40 5.07 69.15 51.58 15.53
PMET 78.60 38.76 81.15 5.26 88.66 49.69 47.58 5.01 53.27 35.73 15.76
ALPHAEDIT 3.03 1.9 54.97 4.21 93.68 51.65 52.33 5.08 80.27 63.66 15.32
NAMET 91.88 61.08 72.01 5.20 95.88 54.08 53.43 5.04 81.24 74.58 15.23
Falcon-7B FT 14.70 13.54 56.34 1.67 23.94 50.46 49.69 3.51 13.64 12.68 32.28
ROME 12.85 12.56 51.48 3.53 74.57 52.10 53.64 5.10 8.39 7.3 10.29
MEMIT 89.21 60.85 77.56 5.19 96.04 55.23 56.91 4.97 82.93 68.93 33.64
PMET 77.61 57.03 70.48 5.17 58.03 54.40 54.49 5.00 69.73 60.69 35.34
ALPHAEDIT 87.62 58.32 72.43 5.00 96.22 55.47 58.02 4.93 53.78 40.83 22.60
NAMET 92.18 63.91 73.86 5.21 96.94 57.08 58.58 5.07 83.18 77.55 36.71
Deepseek-7B FT 2.61 2.49 81.43 5.19 18.85 48.90 52.78 5.00 15.00 12.28 39.14
ROME 0.26 0.30 49.82 5.14 0.55 43.17 46.02 4.06 0.81 0.78 0.75
MEMIT 62.11 42.01 78.04 5.12 33.65 52.28 49.05 4.99 57.10 42.58 39.12
PMET 74.52 43.49 79.01 5.14 86.75 57.85 59.93 5.00 76.97 69.22 39.87
ALPHAEDIT 22.51 14.00 59.92 4.79 18.53 48.33 48.74 4.83 73.41 57.09 34.87
NAMET 89.74 59.98 77.73 5.13 95.14 57.43 60.45 4.97 77.62 71.02 39.16
LLaMA3-8B FT 2.68 1.30 58.16 4.34 16.05 47.51 48.84 4.90 11.75 10.48 42.53
ROME 51.01 33.32 64.37 4.91 82.19 54.94 57.74 5.18 7.40 6.82 27.79
MEMIT 93.76 61.98 77.69 5.26 92.63 55.60 58.75 5.27 78.40 71.76 39.21
PMET 77.71 49.41 71.43 5.10 75.81 56.89 58.26 5.13 68.52 62.72 39.35
ALPHAEDIT 58.97 33.02 85.16 5.37 65.36 51.44 53.55 5.16 64.01 57.01 40.82
NAMET 93.87 65.74 75.07 5.33 93.82 57.88 59.48 5.28 79.34 74.73 44.39
We demonstrate the effectiveness of NAMET in massive editing tasks by comparing it with baseline methods across six popular LLMs. Specifically, we simultaneously edit 10,000 factual associations sampled from the CounterFact and ZsRE datasets. For the Wiki-recent dataset, we modify all 1,266 knowledge items. Additional experiments on other LLMs and datasets are provided in Appendix D. As shown in Table 3, our key findings are as follows:
Table 4: Performance comparison of different editing methods on Qwen2.5-7B, Falcon-7B, and LLaMA3-8B with 15,000 edits from the CounterFact benchmark.
Model Method Counterfact (15000)
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Flu.$\uparrow$
Qwen2.5-7B MEMIT 77.46 54.34 66.23 5.14
NAMET 83.66 55.31 69.49 5.28
Falcon-7B MEMIT 84.60 56.13 75.51 5.13
NAMET 89.55 61.00 68.44 5.16
LLaMA3-8B MEMIT 87.58 54.76 72.65 5.14
NAMET 91.22 62.24 73.43 5.31
Finding 2. NAMET consistently achieves superior editing performance across all datasets and model architectures. Across all evaluated datasets, NAMET demonstrates the highest levels of editing efficacy and generalization. On the CounterFact dataset, it consistently outperforms other methods, particularly on base models such as LLaMA2-7B, LLaMA2-13B, and Deepseek-7B. For example, NAMET achieves 89.09% efficacy and 61.21% generalization on LLaMA2-7B, outperforming the second-best method (PMET) by 15% on both metrics. The gap widens further compared to MEMIT, with improvements of 65% in efficacy and 39% in generalization. Even on more advanced models such as Qwen2.5-7B, Falcon-7B, and LLaMA3-8B, NAMET consistently surpasses all baselines. Furthermore, its advantage becomes more pronounced at larger editing scales. As shown in Table 4, when editing 15,000 knowledge items on Qwen2.5-7B, NAMET achieves 83.66% efficacy, demonstrating a 10% improvement over MEMIT.
Finding 3. NAMET preserves the general abilities of the edited models. In addition to achieving state-of-the-art editing performance, NAMET does not impair the base model’s fluency or reasoning abilities. Across all datasets, NAMET consistently attains among the highest specificity and fluency scores. Notably, on the Wiki-recent dataset, NAMET achieves the best portability performance on most base models, indicating that the edited models retain their ability to reason about downstream knowledge related to the edited facts.
6.4 Robustness against long prefixes (RQ3)
Figure 6: Editing performance of different methods across varying prefix lengths.
We evaluate the robustness of editing methods when edited facts are preceded by varying numbers of tokens. Specifically, we modify 200 facts from the CounterFact dataset in LLaMA2-7B, Deepseek-7B, and Qwen2.5-7B. During evaluation, we prepend prefixes of 5, 50, 100, and 200 tokens to them.
Finding 4. NAMET remains effective when edits are preceded by long prefixes. In Figure 6, NAMET consistently achieves the highest editing efficacy across all models. The maximum efficacy drop with 200-token prefixes is only 7% across the three models. In contrast, MEMIT’s efficacy decreases significantly from 84.75% to 66.50% on LLaMA2-7B, and from 94.2% to 82.25% on Deepseek-7B as prefix length increases. PMET also shows weak robustness, with efficacy drops of 17% and 30% on LLaMA2-7B and Deepseek-7B, respectively. Notably, all methods demonstrate strong robustness on Qwen2.5-7B, consistent with our earlier observation that Qwen2.5-7B is more suitable for robust batch editing. Nevertheless, NAMET exhibits the smallest efficacy drop (only 1.9%) when the prefix increases to 200 tokens, which is half that of the second-best method (MEMIT).
7 Limitations and Future Discussions
While NAMET demonstrates strong performance in massive knowledge editing, its potential in sequential knowledge editing remains under-explored. Specifically, recent work [6] on preserving model performance during sequential editing suggests that the hidden representations of pre-edited and post-edited models should remain similar. This contrasts with our findings, which show that more dispersed hidden representations enable the concurrent editing of a larger number of knowledge items. Future work should explore the balance between sequential and massive editing, as well as between uniform and dispersed hidden representations, in order to facilitate the application of model editing in real-world scenarios.
8 Conclusion
In this paper, we propose NAMET, a novel model editing method that enables stronger and more robust massive editing across various models and datasets. We first identify that the failures of existing methods in massive editing, particularly when evaluated with stricter metrics and context-rich settings, stem from conflicts between residual-key (r-k) and key-key (k-k) pairs. NAMET addresses this issue with a simple modification to MEMIT, introducing noise during the optimization of memory representations to promote more isolated r-k pairs, thereby reducing collisions. Extensive experiments on multiple base LLMs, including LLaMA2, LLaMA3, Deepseek, and Qwen, demonstrate that NAMET significantly outperforms existing methods in editing performance.
Appendix A Proof of Theorem 1
A.1 Problem Setting
We consider the following linear inverse problem that arises during massive editing in MLP layers. Suppose we aim to edit $N_{t}$ knowledge items to a pre-trained LLM. Each edit is associated with:
  • A key vector $k_{i}\in\mathbb{R}^{n}$, which encodes the the factual subject;
  • A residual vector $r_{i}\in\mathbb{R}^{m}$, representing the target object.
Let:
$S_{i}:=k_{i}k_{i}^{T}\in\mathbb{R}^{n\times n},\quad T_{i}:=r_{i}k_{i}^{T}\in\mathbb{R}^{m\times n}.$ (12)
That is, each $S_{i}$ is the outer product of the key with itself (key-key embedding), and $T_{i}$ is the outer product of residual and key (residual-key embedding). The update on MLP layers aims to solve the following equation:
$\Delta\left(C_{p}+\sum_{i=1}^{N_{t}}S_{i}\right)=\sum_{i=1}^{N_{t}}T_{i},$ (13)
where $C_{p}$ denotes the preserved key embeddings that should remain unaffected by the update. This leads to the additional constraint $\Delta C_{p}=0$, which transforms Equation 13 into:
$\Delta S=T,\quad\text{where}\quad S:=\sum_{i=1}^{N_{t}}S_{i},\quad T:=\sum_{i=1}^{N_{t}}T_{i}.$ (14)
We consider the minimum-norm solution for the update:
$\Delta:=TS^{+},$ (15)
where $S^{+}$ denotes the Moore-Penrose pseudoinverse of $S$. The reconstructed target for item $i$ is then:
$\hat{T}_{i}:=\Delta S_{i},$ (16)
and the reconstruction loss is then:
$RL_{i}:=\hat{T}_{i}-T_{i}=\Delta S_{i}-T_{i}.$ (17)
A.2 Error Decomposition
We begin by expanding:
$\Delta=\left(\sum_{j}T_{j}\right)S^{+}\quad\Rightarrow\quad RL_{i}=\left(\sum_{j}T_{j}\right)S^{+}S_{i}-T_{i}.$ (18)
This can be decomposed as:
$RL_{i}=(T_{i}S^{+}S_{i}-T_{i})+\sum_{j\neq i}T_{j}S^{+}S_{i},$ (19)
where:
  • The first term, $T_{i}S^{+}S_{i}-T_{i}$, is the self-alignment bias, reflecting how well $T_{i}$ is reconstructed through its own $S_{i}$;
  • The second term, $\sum_{j\neq i}T_{j}S^{+}S_{i}$, is the interference error, reflecting how other items’ r-k embeddings $T_{j}$ leak into $\hat{T}_{i}$ due to overlap in k-k embeddings.
A.3 Impact of Non-Orthogonality among $\{S_{i}\}$ (Key-Key Collisions)
Assume $S$ is full-rank and well-conditioned, so we approximate:
$S^{+}\approx S^{T}(SS^{T})^{+}.$ (20)
Then:
$T_{j}S^{+}S_{i}\approx T_{j}S^{T}(SS^{T})^{+}S_{i}.$ (21)
This term is large when $S_{i}$ shares directions with other $S_{j}$ (i.e., $S_{i}$ is not orthogonal to $S_{j}$). More precisely, we can express:
$S^{+}S_{i}=\sum_{k}w_{ik}S_{k}^{T},\quad\Rightarrow\quad T_{j}S^{+}S_{i}=\sum_{k}w_{ik}T_{j}S_{k}^{T}.$ (22)
Thus, if $S_{i}$ is not orthogonal to other $S_{k}$, then $S_{i}$ shares eigen-directions with $S_{k}$, and $\hat{T}_{i}$ becomes vulnerable to contamination from $T_{j}$.
A.4 Impact of Non-Orthogonality among $\{T_{i}\}$ (Residual-Key Collisions)
Now suppose r-k embeddings are redundant across edits:
$T_{i}=\bar{T}_{i}+\sum_{j\neq i}\gamma_{ij}T_{j}.$ (23)
where $\bar{T}_{i}$ reflects the real r-k embedding of target item $i$. Then:
$T=\bar{T}_{i}+\sum_{j\neq i}(1+\gamma_{ij})T_{j},\quad\Rightarrow\quad\Delta=TS^{+}=\bar{T}_{i}S^{+}+\sum_{j\neq i}(1+\gamma_{ij})T_{j}S^{+}.$ (24)
Hence,
$\hat{T}_{i}=\Delta S_{i}=\bar{T}_{i}S^{+}S_{i}+\sum_{j\neq i}(1+\gamma_{ij})T_{j}S^{+}S_{i}.$ (25)
The interference term is scaled both by k-k overlap (through $S^{+}S_{i}$) and by r-k redundancy (via $\gamma_{ij}$).
A.5 Conclusion: Double Embedding Collision Amplifies Error
We restate the main result of this section as Theorem 1.
Theorem 1. Given keys $k_{i}\in\mathbb{R}^{n}$ and residuals $r_{i}\in\mathbb{R}^{m}$, define $S_{i}=k_{i}k_{i}^{T}$ and $T_{i}=r_{i}k_{i}^{T}$. Let $S=\sum_{i}S_{i}$, $T=\sum_{i}T_{i}$, and $\Delta=TS^{+}$. Then the reconstruction error $RL_{i}:=\Delta S_{i}-T_{i}$ can be decomposed as:
$RL_{i}=(T_{i}S^{+}S_{i}-T_{i})+\sum_{j\neq i}T_{j}S^{+}S_{i}.$ (26)
The magnitude of $RL_{i}$ is correlated with both:
$\sum_{j\neq i}\cos(k_{i}k_{i}^{T},k_{j}k_{j}^{T})\quad\text{and}\quad\sum_{j\neq i}\cos(r_{i}k_{i}^{T},r_{j}k_{j}^{T}),$ (27)
which we refer to as the embedding collision scores in the key-key and residual-key embeddings, respectively. These collisions hinder the disentanglement of individual edits in the shared parameter space, thereby degrading editing performance under massive updates.
Appendix B Detailed Experiment Setup
In the following, we provide detailed experimental configurations, including the description of the datasets, introduction of baselines, explanation of evaluation metrics, and implementation details.
B.1 Datasets and Baselines
We evaluate the performance of model editing techniques using the following datasets:
  • CounterFact [19] is a benchmark for evaluating factual knowledge localization and editing in LLMs. It contains 21,917 entries that describe the named entities along with their counterfactual variations. Model editing techniques could be evaluated in terms of editing efficacy, generalization, and locality. The benchmark also contains generation prompts to test the model’s generation ability after editing.
  • ZsRE [14] is a question-answering (QA) benchmark designed to evaluate zero-shot relation extraction capabilities of language models. Entries in the benchmark consist of a subject entity along with an answer as the editing target. The benchmark also includes paraphrased questions for testing generalization ability and irrelevant questions for evaluating the locality of editing techniques.
  • Wiki-recent [37] contains 1,266 entries of triplets that have been added into WikiData after July 2022. The benchmark enables insertion for models that were trained prior to the introduction of these facts. This simulates the cases of editing outdated models with newly introduced facts. Model editing techniques are evaluated in terms of editing efficacy, portability, and locality. Here, portability emphasizes whether the edited model could reason about the downstream effects of facts when they are inserted into the model.
We proceed to introduce baseline methods evaluated in the paper. For all baseline methods, we use the official implementation provided by the authors.
  • MEND [21] requires extra parameters for efficiently editing pretrained LLMs. It introduces a set of small auxiliary networks that transform standard fine-tuning gradients into low-rank updates, enabling fast and localized edits without retraining the entire model. This approach offers a scalable solution for post-hoc model editing, avoiding the overfitting issue of traditional fine-tuning methods.
  • ROME [19] performs factual knowledge editing by directly modifying the feed-forward weights in specific layers of LLMs. It first identifies that factual knowledge is primarily stored in mid-layer feed-forward modules, thereby demonstrating the feasibility of editing model parameters to update internal knowledge. The method then updates these weights to encode specific factual associations. ROME achieves precise insertion of new facts with minimal interference to unrelated knowledge.
  • MEMIT [20] is designed to efficiently update LLMs with thousands of factual associations simultaneously. Building upon ROME, MEMIT employs a least-squares optimization over multiple key-value memory components, ensuring high specificity and minimal interference with unrelated knowledge. It further distributes the updates across multiple layers, which helps reduce the impact on the model’s general capabilities.
  • PMET [15] is a method designed to enhance the precision of knowledge updates in large language models. Unlike prior approaches that treat transformer layer (TL) hidden states as direct inputs of the feed-forward network (FFN), PMET recognizes that these hidden states also encompass information from multi-head self-attention (MHSA) and residual connections. PMET proceeds to simultaneously optimize MHSA and FFN hidden states and use the optimized TC hidden states of FFN to precisely update FFN weights. This approach enables more accurate and efficient model editing, preserving the integrity of the model’s existing knowledge while incorporating new information.
  • ALPHAEDIT [6] preserves knowledge in LLMs during sequential updates by projecting updates onto the null space of the preserved knowledge. This could ensure that new modifications do not interfere with previously stored information. This approach maintains the integrity of the model’s existing knowledge while enabling precise edits.
B.2 Metrics
We now introduce the metrics used for CounterFact, Wiki-recent and ZsRE respectively.
B.2.1 CounterFact Metrics
Given an LLM $f_{\theta}$, a knowledge fact tuple (subject $s_{i}$, relation $r_{i})$, a target output $o_{i}$ and the original output $o_{i}^{c}$, we define the following metrics:
  • Editing Efficacy: Unlike previous works that evaluate the portion of cases where $o_{i}$ is more probable than $o_{i}^{c}$, we directly compute the average top-1 accuracy of edited samples.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ (s_{i},r_{i}))]$ (28)
  • Generalization: Average top-1 accuracy of the edited model on rephrased statements $N((s_{i},r_{i}))$ of the original knowledge fact. Rephrased statements share the same semantic meaning with the original statements.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ N((s_{i},r_{i})))]$ (29)
  • Specificity: The portion of cases where $o_{i}^{c}$ is more probable than $o_{i}$ with neighboring statements $O((s_{i},r_{i}))$. Neighboring statements are constructed using prompts which share distinct but semantically related subjects with the original knowledge fact.
    $\mathds{E}_{i}[\mathds{P}_{f_{\theta}}(o_{i}^{c}\ |\ O((s_{i},r_{i})))>\mathds{P}_{f_{\theta}}(o_{i}\ |\ O((s_{i},r_{i})))]$ (30)
  • Fluency: Fluency score measures the quality of the generated text. It scores low if the generated text contains excessive repetition.
    $-\frac{2}{3}\mathop{\sum}_{k}g_{2}(k)\log_{2}g_{2}(k)+\frac{4}{3}\mathop{\sum}_{k}g_{3}(k)\log_{2}g_{3}(k)$ (31)
    where $g_{2}(k)$ and $g_{3}(k)$ are the probabilities of bigram and trigram $k$ respectively.
B.2.2 Wiki-recent Metrics
Given a LLM $f_{\theta}$, a knowledge fact tuple $(s_{i},\ r_{i})$, a target output $o_{i}$ and the original output $o_{i}^{c}$, we define the following metrics:
  • Editing Efficacy: Average top-1 accuracy of edited samples.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ (s_{i},r_{i}))]$ (32)
  • Portability: Average top-1 accuracy of the edited model on portability prompts $P((s_{i},r_{i}))$ of the original knowledge fact. Portability prompts contain three parts: alias prompts, compositionality and reasoning prompts, and logical generation prompts. Specifically, alias prompts are constructed by replacing the subject $s_{i}$ with an alias or synonym. Compositionality and reasoning prompts require the post-edit model to conduct reasoning about the changed fact. Logical generation prompts are changes that are semantically related to the modified fact and expected to change by the edit.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ P((s_{i},r_{i})))]$ (33)
  • Locality: Average top-1 accuracy of the edited model on neighboring prompts $O((s_{i},r_{i}))$ of the original knowledge fact.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ O((s_{i},r_{i})))]$ (34)
  • Fluency: Fluency score measures the quality of the generated text. It scores low if the generated text contains excessive repetition.
    $-\frac{2}{3}\mathop{\sum}_{k}g_{2}(k)\log_{2}g_{2}(k)+\frac{4}{3}\mathop{\sum}_{k}g_{3}(k)\log_{2}g_{3}(k)$ (35)
    where $g_{2}(k)$ and $g_{3}(k)$ are the probabilities of bigram and trigram $k$ respectively.
B.2.3 ZsRE Metrics
Given a LLM $f_{\theta}$, a knowledge fact tuple $(s_{i},\ r_{i})$, a target output $o_{i}$ and the original output $o_{i}^{c}$, we define the following metrics:
  • Editing Efficacy: Average top-1 accuracy of edited samples.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ (s_{i},r_{i}))]$ (36)
  • Generalization: Average top-1 accuracy of the edited model on generalization prompts $N((s_{i},r_{i}))$ of the original knowledge fact.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ N((s_{i},r_{i})))]$ (37)
  • Locality: Average top-1 accuracy of the edited model on neighboring prompts $O((s_{i},r_{i}))$ of the original knowledge fact.
    $\mathds{E}_{i}[o_{i}=\arg\mathop{\max}_{o}\mathds{P}_{f_{\theta}}(o\ |\ O((s_{i},r_{i})))]$ (38)
B.2.4 Examples of Evaluation Prompts with Prefixes
We provide examples of evaluation prompts with prefixes in the following. For the evaluated prompt "The mother tongue of Danielle Darrieux is French", corresponding evaluated prompts with ten distinct 5-token prefixes are:
Example Prompts with Prefixes The doctor’s office was too. The mother tongue of Danielle Darrieux is French. Therefore, the average speed of. The mother tongue of Danielle Darrieux is French. You can find many examples of. The mother tongue of Danielle Darrieux is French. However, the government’s new. The mother tongue of Danielle Darrieux is French. And so, as the night. The mother tongue of Danielle Darrieux is French. While the world of sports and. The mother tongue of Danielle Darrieux is French. To make a cake, you. The mother tongue of Danielle Darrieux is French. Nevertheless, the overall sentiment of. The mother tongue of Danielle Darrieux is French. Never tried it before. The mother tongue of Danielle Darrieux is French. He realized that the people of. The mother tongue of Danielle Darrieux is French.
For model evaluation, we generate prefixes by prompting the unedited model with the following initial words: "The", "Therefore", "You", "However", "And", "While", "To", "Nevertheless", "Never", and "He". These initial words are used to generate diverse 5-token prefixes, which are then prepended to each edited fact during the evaluation process. This approach ensures a comprehensive assessment of the model’s performance across different linguistic contexts.
B.3 Implementation Details
We implement all experiments on a single NVIDIA H800 GPU with 80GB memory. The details of our implementation across different models are outlined as follows:
  • LLAMA2-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [25000, 50000] respectively for two layers. During optimization, we iterate for 25 steps with 0.04 learning rate.
  • Qwen-2.5-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [500, 1000] respectively for two layers. During optimization, we iterate for 25 steps with 0.04 learning rate.
  • LLaMA2-13B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [25000, 50000] respectively for two layers. During optimization, we iterate for 25 steps with 0.04 learning rate.
  • Falcon-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [5000, 1000] respectively for two layers. During optimization, we iterate for 25 steps with 0.06 learning rate.
  • Deepseek-base-7B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [10000, 20000] respectively for two layers. During optimization, we iterate for 25 steps with 0.06 learning rate.
  • LLaMA3-8B: We modify layers [3, 4] for editing factual knowledge. The hyperparameters $\lambda$s are set to [7500, 15000] respectively for two layers. During optimization, we iterate for 25 steps with 0.008 learning rate.
Appendix C Algorithmic Description of NAMET
In this section, we provide a detailed algorithmic description of NAMET (Figure 4). Following previous work, we also spread MLP updates evenly across target critical layers $\mathcal{R}$. We define $L$ as the final layer of candidate layers, where new memories are fully represented. For each target edit $(s_{i},r_{i},o_{i})$, we first compute the target hidden representation $z_{i}$ to replace the original representation $h_{i}^{L}$ for modifying new factual associations. During the optimization procedure, the padding strategy is changed to left-padding to introduce additional noise for more isolated target representations. We proceed to separately modify MLP modules at every critical layer $l\in\mathcal{R}$ with evenly distributed change $\delta_{i}=z_{i}-h_{i}^{L}$ for every target edit $i$.
Optimizing target memory $z_{i}$. For each subject-relation pair $(s_{i},r_{i})$, we concatenate them into prompt $tp(s_{i},r_{i})$. We proceed to augment each prompt with additional $N_{FP}-1$ random prefixes of fixed length. The original prompt, together with its augmented companions, is tokenized using a left-padding tokenizer, introducing [unk] tokens to the beginning of the original prompt (as shown in Figure 4). These [unk] tokens carry no semantic meaning and provide additional noise during optimization. The target representation $z_{i}$ is then optimized via:
$z_{i}=h_{i}^{L}+\underset{\delta_{i}}{\arg\min}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log G_{(h_{i}^{L}+=\delta_{i})}[o_{i}^{t}\mid f_{j}\oplus tp(s_{i}^{t},r_{i}^{t})]$ (39)
where $\{f_{j}\oplus tp(s_{i},r_{i})\}$ concatenates random prefixes $f_{j}$ with templated prompt $tp(s_{i},r_{i})$ for each target edit $i$. $G_{(h_{i}^{L}+=\delta_{i})}$ denotes that we add change $\delta_{i}$ to the hidden state $h_{i}^{L}$ during the model execution.
Spreading $\delta_{i}$ across candidate layers. Since the editing representation at earlier layers affects that of subsequent layers, we perform edits in ascending layer order. For each edit, we distribute the desired change $\delta_{i}$ proportionally across all candidate layers. Specifically, the residual to be added at layer $l\in\mathcal{R}$ is defined as:
$r_{i}^{l}=\frac{z^{i}_{t}-h^{L}_{i}}{L-l+1}$ (40)
where $h^{L}_{i}$ denotes the hidden state at the final layer $L$ before editing layer $l$.
In addition to $r_{i}^{l}$, we require the corresponding key $k_{i}^{l}$ to update the MLP module. For each subject $s_{i}$, the key is computed as:
$k_{i}^{l}=\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}k(f_{j}\oplus s_{i}^{t}),\quad\text{where}\quad k(x)=\sigma(W_{in}^{l}\gamma(h^{l-1}_{i}(x)))$ (41)
Here, $f_{j}$ is the same random prefix concatenated with the subject $s_{i}$ mentioned above. Notably, prepending [unk] tokens does not affect key computation, thereby preserving the accuracy of the extracted key representations. After aggregating keys $K_{t}=[k_{1}^{t}\ |\ k_{2}^{t}\ |\ ...\ |\ k_{N_{t}}^{t}]$ and residuals $R_{t}=[r_{1}^{t}\ |\ r_{2}^{t}\ |\ ...\ |\ r_{N_{t}}^{t}]$ of all $N_{t}$ target edits, individual $\Delta^{l}$ for candidate layers could then be computed using Equation 4. We summarize the full algorithm in Algorithm C.1.
We provide an intuitive analysis of the effect of introducing [unk] tokens during optimization by examining the norm of hidden states at layer 4, as shown in Table 5. The results indicate that incorporating [unk] tokens introduces additional noise into the hidden states, increasing their norm by at least 4$\times$. Consequently, optimizing over these noisier representations can lead to more isolated and disentangled target representations.
Table 5: Hidden state norm of NAME and MEMIT across different models.
Model NAMET MEMIT
LLaMA2-7B 1321.3±114.6 6.7±0.9
Qwen-2.5-7B 21818.1±2399.9 45.7±5.7
LLaMA2-13B 670.5±26.5 7.3±0.9
Falcon-7B 257.4±2.9 32.0±3.4
Deepseek-7B 6275.6±192.9 23.9±2.8
LLaMA3-8B 570.2±25.4 3.7±0.6
Gemma-7B 70.9±6.1 15.6±2.8
Phi-1.5 209.5±14.2 48.3±5.6
1
2Data: Requested edits =$\mathcal{E}=\{(s_{i},r_{i},o_{i})\}$, generator $G$, layers to edit $\mathcal{S}$, covariances $C^{l}$
3 Result: Modified generator containing edits from $\mathcal{E}$
4
5 tokenizer.padding_side $\leftarrow$ "left"
6  // Change padding strategy to left
7 for $s_{i},r_{i},o_{i}\in\mathcal{E}$ do
8       optimize $\delta_{i}\leftarrow\arg\min_{\delta_{i}}\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}-\log\mathbb{P}_{G_{(h_{i}^{L}+=\delta_{i})}}[o_{i}\mid f_{j}\oplus tp(s_{i},r_{i})]$
9       $z_{i}\leftarrow h_{i}^{L}+\delta_{i}$
10
11
12for $l\in\mathcal{R}$ do
13       $h_{i}^{l}\leftarrow h_{i}^{l-1}+a_{i}^{l}+m_{i}^{l}$
14       for $s_{i},r_{i},o_{i}\in\mathcal{E}$ do
15             $k_{i}^{l}\leftarrow k_{i}^{l}=\frac{1}{N_{FP}}\sum_{j=1}^{N_{FP}}k(f_{j}\oplus s_{i})$
16             $r_{i}^{l}\leftarrow\frac{z_{i}-h_{i}^{L}}{L-l+1}$
17              // Distribute over remaining layers
18
19      $K^{l}\leftarrow[k_{1}^{l},\ldots,k_{N_{t}}^{l}]$
20       $R^{l}\leftarrow[r_{1}^{l},\ldots,r_{N_{t}}^{l}]$
21       $\Delta^{l}\leftarrow R^{l}K^{l^{T}}(C^{l}+K^{l}K^{l^{T}})^{-1}$
22       $W^{l}\leftarrow W^{l}+\Delta^{l}$
23        // Update layer $l$ MLP weights in model
24
Algorithm C.1 The NAMET Algorithm
Appendix D Additional Experimental Results
In this section, we present additional experimental results and findings to further demonstrate the effectiveness of NAMET. First, we evaluate editing performance across different semantic categories. Next, we show that NAMET minimally interferes with the model’s general capabilities by assessing performance on the GLUE benchmarks. We then report editing performance on two additional LLMs: Gemma-7B[29] and Phi-1.5[16]. Finally, we provide a comprehensive analysis of the hyperparameters.
D.1 Editing Performance Involving Different Semantics
Figure 7: Performance comparison of different editing methods across different semantics.
Additional Finding 1. NAMET achieves superior editing performance across different semantics. We extract samples with specific relation types from the CounterFact dataset to evaluate the performance of different editing methods across semantic categories. As shown in Figure 7, NAMET consistently achieves the highest editing efficacy and generalization on both LLaMA2-7B and Qwen2.5 for most semantic types. On LLaMA2-7B, NAMET outperforms the second-best method (PMET) by approximately 10% in efficacy and 20% in generalization. On the “twin city” relation, NAMET achieves twice the efficacy and four times the generalization of PMET. MEMIT, in contrast, nearly fails to edit facts with this relation, with both efficacy and generalization scores close to 0%. In terms of editing specificity, NAMET performs better on Qwen2.5 than on LLaMA2. On LLaMA2, PMET surpasses NAMET by an average of 5%, whereas on Qwen2.5, NAMET achieves the highest specificity on 6 out of 8 relation types.
D.2 General Ability of Edited Models on GLUE Benchmarks
Figure 8: General ability of pre-edited model and models edited by different methodson GLUE benchmarks
Additional Finding 2. NAMET better preserves the general ability of LLMs after massive editing. We assess whether massive editing degrades the general capabilities of LLMs when using MEMIT, PMET, and NAMET. To this end, we evaluate the performance of three LLMs (LLaMA2-7B, Qwen2.5-7B, and LLaMA3-8B) on the GLUE benchmark after editing 10,000 knowledge facts. For reference, we also report the performance of each unedited (pre-edit) model. As shown in Figure 8, models edited with NAMET exhibit the smallest performance deviation from their pre-edited counterparts. On LLaMA3-8B, the average deviation across six GLUE metrics for the NAMET-edited model is 0.118, which is 0.04 lower than MEMIT and half that of PMET. On LLaMA2-7B, NAMET maintains deviations below 0.05 for 4 out of 6 metrics, with a maximum deviation of only 0.071. The overall average deviation is 0.036, which is the lowest among all methods. NAMET demonstrates a strong ability to preserve general capability on Qwen2.5-7B, with an average deviation of 0.189, significantly lower than those of MEMIT (0.269) and PMET (0.308).
We attribute this robustness to NAMET’s ability to extract more dispersed memory representations across knowledge items. Such dispersion reduces the likelihood of embedding collisions during massive editing, which may compromise the model’s general capabilities. By mitigating this interference, NAMET effectively preserves the original functionality of the LLM.
D.3 Analysis of Editing Performance across LLMs
Table 6: Performance comparison of different editing methods on Gemma-7B and Phi-1.5 on the Counterfact and ZsRE benchmarks.
Model Method Counterfact ZsRE
Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$ Flu.$\uparrow$ Eff.$\uparrow$ Gen.$\uparrow$ Spe.$\uparrow$
Gemma-7B MEMIT 93.01 54.33 74.88 5.38 84.61 73.66 23.04
PMET 91.69 47.24 75.22 5.33 83.05 73.95 23.85
NAMET 95.29 68.53 66.22 5.30 85.83 78.23 23.37
Phi-1.5 MEMIT 49.00 29.67 63.71 5.68 49.76 38.54 20.36
PMET 38.49 20.75 67.01 5.79 32.00 24.47 21.54
NAMET 67.76 40.13 62.08 5.80 51.54 41.02 21.29
We further evaluate the performance of NAMET on two additional LLMs: Gemma-7B and Phi-1.5. As shown in Table 6, NAMET consistently achieves the highest editing efficacy and generalization across both models and datasets. Moreover, it maintains competitive performance in terms of editing locality and generation ability.
Table 7: Model Architecture and Editing Efficacy Comparison
Model Hid. size Inter. size Vocab size Eff. (MEMIT) Eff. (NAMET)
LLaMA3-8B 4096 14336 128256 93.76 93.87
Gemma-7B 3072 24576 256000 93.01 95.29
Qwen-2.5-7B 3584 18944 152064 90.06 90.49
Falcon-7B 4544 18176 65024 89.21 92.18
Deepseek-7B 4096 11008 102400 62.11 89.74
Phi-1.5-1.3B 2048 8192 51200 49.01 67.76
LLaMA2-13B 5120 13824 32000 47.98 91.88
LLaMA2-7B 4096 11008 32000 24.95 89.09
We further analyze architectural differences across the evaluated LLMs, focusing on the hidden size, the intermediate size of FFN modules, and the vocabulary size. We report their corresponding editing efficacy when editing 10,000 facts from the CounterFact dataset using both MEMIT and NAMET, as shown in Table 7.
Additional Finding 3. LLMs with larger hidden or intermediate sizes generally demonstrate better tolerance to massive editing. Table 7 shows that editing efficacy under MEMIT tends to increase with the model’s hidden or intermediate size. For example, compared to LLaMA2-7B, LLaMA3-8B has approximately 3,000 more intermediate dimensions and achieves 69% higher editing efficacy. Although Gemma-7B and Qwen2.5-7B have smaller hidden sizes, their intermediate sizes are approximately 1.7$\times$ and 2$\times$ larger than that of LLaMA2-7B, respectively, resulting in 68% and 65% higher efficacy. We attribute this improvement to the increased dimensionality in both the key (intermediate state) and memory (hidden state) representations, which provide more capacity for encoding diverse knowledge embeddings. This increased capacity enables the formation of more dispersed k-k and r-k embeddings across knowledge items, thereby reducing the risk of collisions during editing.
Additional Finding 4. Enlarging the vocabulary size can compensate for smaller hidden or intermediate dimensions. Two notable exceptions to the trend discussed above are Deepseek-7B and Phi-1.5-1.3B, which have hidden and intermediate sizes that are equal to or smaller than those of LLaMA2-7B, yet exhibit higher editing efficacy. We hypothesize that this improvement arises from their significantly larger vocabulary sizes. A larger vocabulary allows the model to learn more fine-grained and distinct token embeddings, which can help differentiate knowledge representations and reduce the likelihood of embedding collisions during editing.
Additional Finding 5. NAMET improves editing efficacy across architectures, particularly for models with smaller hidden or intermediate sizes. As shown in Table 7, NAMET consistently outperforms MEMIT in editing efficacy across all evaluated models. Notably, for models with relatively small hidden or intermediate dimensions, NAMET achieves editing efficacy comparable to that of larger models. For instance, NAMET attains 89.09% editing efficacy on LLaMA2-7B, closely matching the performance of LLaMA3-8B (93.87%), Qwen-2.5-7B (90.49%), and Falcon-7B (92.18%). This improvement is attributed to the introduction of padding tokens, which act as a form of noise during the optimization of memory embeddings. The injected noise encourages the model to learn more diverse and disentangled representations for different knowledge items, thereby reducing the risk of embedding collisions even in models with limited representational capacity.
D.4 Detailed Hyperparameter Analysis
We conduct extensive experiments to analyze the impact of different hyperparameters on the performance of NAMET.
Figure 9: Performance comparison of MEMIT and NAMET across different layers of LLaMA2-7B when editing 100 and 500 knowledge items.
Additional finding 6. Knowledge can be edited at arbitrary layers, but editing at deeper layers is more prone to conflict under massive editing. We modify 100 and 500 knowledge items from the CounterFact dataset in LLaMA2-7B and vary the editing layer from 2 to 30. As shown in Figure 9, editing efficacy remains largely unaffected by the choice of editing layer when editing 100 items. Both NAMET and MEMIT achieve over 90% efficacy at nearly all layers, except the last one. However, when editing 500 items, MEMIT’s editing efficacy gradually declines and drops to only 12.6% at layer 30. This is because r-k pairs extracted from deeper layers carry more information and are thus more likely to collide with others. We further analyze the average $\ell_{2}$ norm of r-k pairs and the number of elements in r-k vectors whose variance exceeds 0.01 across 100 knowledge items, as shown in the right plot of Figure 9. Both metrics show a monotonic increase as the editing layer deepens. Notably, the average $\ell_{2}$ norm of r-k pairs at layer 30 is nearly twice that at layer 28. These results suggest that successful editing at deeper layers requires more isolated r-k representations. NAMET maintains high editing efficacy even at layer 30, achieving 81.2% efficacy.
Figure 10: Impact of varying $\lambda$ and learning rate (lr) on the editing efficacy of NAMET on the CounterFact dataset.
Additional Finding 7. NAMET requires an appropriately tuned $\lambda$ and learning rate to achieve optimal performance. We vary $\lambda$ from 2000 to 50000 and the learning rate from 0.02 to 0.08 to examine their impact on NAMET’s performance when editing 2,000 facts from the CounterFact dataset in LLaMA2-7B. Recall that $\lambda$ scales the k-k embedding $K_{p}K_{p}^{T}$ of preserved knowledge, balancing its influence against that of newly introduced knowledge. As shown in the left plot of Figure 10, the editing efficacy increases and reaches a peak of over 98% as $\lambda$ rises from 2000 to 10000, but then drops to around 90% when $\lambda$ is further increased to 50000. A similar trend is observed in the right plot of Figure 10 for the learning rate: the editing efficacy peaks at a learning rate of 0.04 and deteriorates when the learning rate is either reduced to 0.02 or increased to 0.08.
We further analyze the impact of the regularization coefficient $\lambda$ and the learning rate on the editing efficacy of NAMET. We begin by recalling the final update formulation for the MLP weights:
$\left(C_{p}+\sum_{i=1}^{N_{t}}k_{i}k_{i}^{T}\right)\Delta=\sum_{i=1}^{N_{t}}r_{i}k_{i}^{T},$ (42)
where $C_{p}=\lambda E_{k_{p}}[k_{p}k_{p}^{T}]$ serves as a regularization term derived from preserved knowledge. Each residual $r_{i}$ obtained during editing can be decomposed into two components: $r_{i}=r_{i}^{r}+r_{i}^{n}$, where $r_{i}^{r}$ represents the informative signal and $r_{i}^{n}$ captures the noise introduced by NAMET’s padding tokens.
In this context, $C_{p}$ must not only balance the contribution between preserved and new knowledge, but also help suppress the noise component $r_{i}^{n}$. When $\lambda$ is too small, $C_{p}$ is insufficient to cancel out the noise, resulting in contamination of the residual signal $r^{r}$ and thus reduced editing efficacy. Conversely, an overly large $\lambda$ may over-regularize, suppressing useful information in $r^{r}$ and again degrading performance.
A similar trend exists with respect to the learning rate. If the learning rate is too small, the optimization process fails to effectively encode the target knowledge, resulting in weak or ineffective updates. On the other hand, an excessively large learning rate may amplify the injected noise, distorting the representation and requiring a correspondingly larger $\lambda$ to suppress the resulting artifacts. Furthermore, high learning rates may overwhelm or obscure the true signal in $r_{i}^{r}$, ultimately leading to degraded editing performance.
Appendix E Specific Cases
In this section, we present representative editing examples from the CounterFact dataset to qualitatively assess the generalization performance of different editing methods after editing 10,000 facts on LLaMA2-7B and Qwen2.5-7B. These examples reveal that baseline methods often fail to generate the correct target answers or produce repetitive and incoherent responses. In contrast, NAMET consistently retrieves the intended factual content while producing fluent and informative completions, further validating the robustness and effectiveness of our proposed method.
Model Editing Example on LLaMA2-7B Target Edit The original language of The Icelandic Dream was Original Object Icelandic Target Object Tamil Generation Output NAMET The Icelandic Dream is full of people who speak Tamil, and the film is in Tamil. The film is about a family of Tamil refugees who are trying to make a new life for themselves in Europe. The film is about the struggles of the refugees, and the film is about the struggles of the people who are trying to make a MEMIT The Icelandic Dream is full of people who speak English, and the locals are very friendly. is is, a Belgian chain, has 100 restaurants in the UK, and is opening 10 more this year.,,, a Belgian chain, has 100 restaurants in the UK, and is opening 1 PMET The Icelandic Dream is full of people who speak in a language that is not English. The Icelandic Dream is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the immigrant experience. It is a film about the
Model Editing Example on Qwen2.5-7B Editing Prompt Argentine Football Association belongs to the organization of Original Object FIFA Target Object NATO Generation Output NAMET Argentine Football Association is part of NATO. NATO is military alliance. NATO is also called as what? NATO is a military alliance that was formed in 1949. It is also known as the North Atlantic Treaty Organization. NATO is made up of 29 member countries, including the United States, Canada, and several European countries. MEMIT Argentine Football Association is part of the American Football Association. It is a group of American players who are of African-American descent. The group was formed in 1991 and is based in the United States. The group is a member of the International Football Association. The group is also a member of the American Football Association. The group is also PMET Argentine Football Association is part of NATO’s military alliance with the United States. The alliance is also known as the North Atlantic Treaty Organization or NATO. The alliance is a military alliance between the United States and NATO. The alliance is a military alliance between the United States and NATO. The alliance is a military alliance between the United States and NATO. The alliance [repetitive pattern]